Consider trade in automobiles between the United States and Europe. The average European car costs €15,000. Suppose that theNet Exports NET EXPORTS (Billions of dollars) -1.5 1 2 3 4 9 10 11 12 5 6 7 8 TIME (Month of year)
Answer #1

Net exports (NX) = Exports - Imports

(a) In March,

Total imports = (15,000 x 90,000) x 1.16 = $1.566 billion (= $1,566 million)

NX = 0.52 - 1.566 = - 1.046 billion (= - $1,046 million)

(b) In May,

Total imports = (15,000 x 85,000) x 1.31 = $1.67025 billion (= $1,670.25 million)

NX = 0.52 - 1.67025 = - 1.15025 billion (= - $1,150.25 million)

(c) In October,

Total imports = (15,000 x 75,000) x 1.31 = $1.47375 billion (= $1,473.75 million)

NX = 1.57 - 1.47375 = 0.09625 billion (= $96.25 million)

(d) Graph

Net Exports NET EXPORTS (Billions of dollars) -1.5 - 1 2 3 4 5 6 7 8 TIME (Month of year) 9 10 11 12

NOTE: Since you didn't provide drop-down options, I've also shown values in $Million.